Please type your answer, not hand writing.

Scoresby Inc. tracks the number of units purchased and sold throughout each year but applies Its Inventory costing method atScoresby Inc. tracks the number of units purchased and sold throughout each year but applies Its Inventory costing method atScoresby Inc. tracks the number of units purchased and sold throughout each year but applies Its Inventory costing method atRequired 1 Requires a Reguires 2 Requires Required 2 Required 3 Required 4 Resuires a requires 6 Required 6 Prepare an incomeComplete this question by entering your answers in the tabs below. Required 1 Requires : Requires 2 Requires a Reguires a Reg
Answer #1

1.

Date Unit Unit cost Amount
Beginning inventory 3000 $8 $24,000
March 5 9500 9 85,500
Sep.19 5000 11 55,000
Cost of goods available for sale 17500 $164,500

Average cost per unit = $164,500 / 17,500 = $9.4

Number of goods available for sale = 17,500 units

Cost of goods available for sale = $164,500

2.

Number of units in ending inventory = 17,500 - 12,000 (4,000+8,000) = 5,500 units

3.

Cost of ending inventory Cost of goods sold
FIFO $59,500 ($55,000+500*$9) $105,000 ($164,500-59,500)
LIFO 46,500 (24,000+2500*$9) 118,000 (164,500-46,500)
Weighted average 51,700 (5500*$9.4) 112,800 (12,000*$9.4)

4.

SCORESBY INC.

Income Statement

For The Year Ended December 31

FIFO LIFO Weighted average
Sales (4,000*29+8,000*31) $364,000 $364,000 $364,000
Cost of goods sold 105,000 118,000 112,800
Gross margin 259,000 246,000 251,200
Operating expenses 250,000 250,000 250,000
Income (loss) from operations $9,000 $(4,000) $1,200

5.

LIFO method minimizes income tax.